OpenStudy (kitten_is_back):

Find the mean of this set of numbers. 5, –11, 7, 12, –1, 0 A. 6 B. –6 C. –2 D. 2 a

OpenStudy (kaylardz0405):

add them all and then subtract by the amount of numbers there is

OpenStudy (kaylardz0405):

i mean divide

OpenStudy (anonymous):

The answer is D. You add them all together then divide by the amount of numbers there are.

OpenStudy (kaylardz0405):

yes 2 would be the answer

OpenStudy (texaschic101):

true ^^....when you have negative numbers in there, you just add them in there just like you do positive numbers

OpenStudy (texaschic101):

5, -11, 7, 12, -1,0 = 24 - 12 = 12 now divide by 6 because there are 6 numbers 12/6 = 2